# DepthCues: Evaluating Monocular Depth Perception in Large Vision Models

### v1

@@ 1 Introduction
Large-scale pre-trained vision models like CLIP [57], Stable Diffusion [61], and DINOv2 [55], trained solely on 2D images, have shown strong generalization when fine-tuned on various computer vision tasks, including monocular depth estimation [17, 22, 81], even reaching human-level performance in ordering objects by depth [47]. As a result, numerous recent monocular depth estimation approaches [28, 39, 78] leverage these models’ depth priors to achieve state-of-the-art performance. This naturally raises an important question: How does depth perception arise in these large vision models?
@@ 2 Related Work
Monocular Depth Estimation. Monocular depth estimation is an inherently ill-posed and challenging problem, yet deep learning-based methods have made significant strides in addressing it. Initial progress was driven by supervised methods [16, 19, 6] that were trained on various labeled datasets [53, 23]. Later, self-supervised approaches advanced the field further by leveraging stereo [21, 25] or temporal consistency [85, 27, 26]. More recently, general purpose methods [59, 77, 78, 39] leveraging large vision backbone models pre-trained with self-supervised [55] or generative objectives [61] and then fine-tuned on very large-scale supervised datasets have been shown to result in strong monocular depth estimation performance. However, it remains unclear how depth perception develops in the pre-trained models without explicit depth supervision during pre-training. In this paper, we try to shed light on the emergence of these monocular depth cues used by the human visual system in large vision models.
@@ 3.4 Perspective
Task. This cue is closely tied to vanishing point [48], which is defined as the point on the image plane where 2D perspective projections of 3D parallel lines intersect. Therefore, we use the task of vanishing point estimation to investigate the understanding of perspective. While there exist different versions of the task that define different numbers of vanishing points [67], we adopt a simplified version that focuses on a single dominant vanishing point. Given an image $I$, the objective is to estimate the $\mathbf{p}=(x,y)$ coordinates of the vanishing point in the image coordinate system.
@@ 4.2 Probing Features
Following [55, 81], we perform a hyper-parameter search on the layers of model $\phi$. For each task, we extract features from different layers of $\phi$, and train a probe on these features under the same setting. The best layer is then selected based on the validation performance, and the probing on this layer is repeated five times using different random seeds. We report the mean and standard deviation of the test performance.
@@ 5.1 Monocular Depth Cues in Vision Models
Additional Baselines. During the probing of certain cues, we explicitly use mask information to construct the input feature, $f$. Although these masks are not directly used by the probes, they could theoretically be leveraged to solve tasks. To verify that any additional information from this step is insufficient to solve the task, we designed a trivial baseline in which model features are replaced with simple 2D coordinate maps i.e., $\phi(I)$ in Eqs. 1 and 2 becomes $C\in\mathbb{R}^{H\times W\times 2}$ where $C_{h,w}=[h,w]^{T}$. Moreover, for each cue, we include an end-to-end baseline, where we train ResNet18 [29] and ViT-B/16 [15] models from scratch in an end-to-end fashion and select the best. We concatenate the masks with the input images channel-wise, allowing us to assess how much of each task can be solved without relying on sophisticated pre-trained features. Finally, to confirm that the tasks are readily solvable by humans, we evaluate our own performance on a random sample from the test sets, and obtained an accuracy of 95%$\pm$1.48across tasks.
Additional Tasks. To extend our analysis, we also evaluate vision models for monocular depth estimation (on NYUv2 [63]) and image classification (on ImageNet-1k [14]) using a linear probe, a standard approach in previous work [55, 79, 3] for probing these datasets. Following the previous works [6, 11, 17], For NYUv2, we report accuracy as the percentage of pixels where the ratio between predicted and ground-truth depth is less than 1.25. We report top-1 accuracy for ImageNet. More details on these tasks are provided in Sec. A.1, and we show the linear probing results of these models on NYUv2 in Fig. 3.
@@ 5.3 Discussion
Limitations. Our analysis focuses only on widely used pre-trained model weights but not directly on the impact of various pre-training objectives and their training datasets on the emergence of depth cues. While depth from ego-motion and scene motion are also important cues for humans, our focus is primarily on assessing monocular cues learned by large vision models trained on static image collections and not image sequences, except CroCo, LRM, and DUSt3R.
@@ 6 Conclusions
Acknowledgments.
Funding was provided by ELIAI (the Edinburgh Laboratory for Integrated Artificial Intelligence), EPSRC (grant no. EP/W002876/1).
@@ A.1 Full Benchmark Results
The probing results of all models are summarized in Tab. A1, and the average performance of models on DepthCues is plotted against their NYUv2 depth estimation results in Fig. A1. It is noted that there are slight differences between the linear probing results on NYUv2 and ImageNet-1k compared to previous works [55, 17]. These may stem from differences in the implementation details including the features probed, learning rates, and training iterations, etc. For example, the probing results on NYUv2 depth estimation in [17] are generally higher than our results. Possible reasons for this include the following factors. Firstly, while we train a simple linear layer to probe the patch tokens from the last layer of the vision models, a more complex non-linear probe similar to the DPT decoder [60] is used to probe features from multiple layers of the models in [17]. Secondly, the class tokens are utilized for several models (e.g., DINO, DINOv2) in [17], while we only use patch tokens for consistent evaluation across all models as some of the models do not include a class token. Thirdly, because of the difference in the probes, their optimization hyper-parameters can differ. Similar reasons can be applied to the differences in our ImageNet-1k probing results from previous works. In particular, for CLIP, the class token seems to be crucial for its ImageNet classification ability. We made these design choices because our primary focus is to evaluate the awareness of human-like monocular depth cues in vision models in a fair and comparable manner, rather than optimizing their probing performance on external benchmarks.
@@ A.5 Learning Depth Cues
In Sec. 5.2 of the main paper, we presented linear probing results on NYUv2 and DIW, showing that fine-tuning on DepthCues improves models’ performance on depth estimation. Here we further investigate whether the fine-tuned model has improved understanding of the depth cues by comparing the probing results of the fine-tuned and pre-trained DINOv2 on DepthCues. Note we focus on probing the last layer since LoRA was only applied on that layer during fine-tuning. The results from Tab. A3 shows that the fine-tuned version outperforms the original DINOv2 on four cues, with a slight drop in performance on light-shadow and perspective. Notably, we see a 9.4% increase in accuracy on texture-grad, on which the original DINOv2 shows relatively weaker understanding (see Fig. 3 in the main paper).
@@ C.2 Evaluation Metrics
While we use accuracy to evaluate performance on the binary classification tasks, as discussed in the main paper we apply thresholding to the Euclidean distance (for perspective) and horizon detection error [^2] This measures the maximum distance between the predicted and the ground-truth horizon lines, normalized by the image height. [75] (for elevation) to convert these metrics into accuracy. Specifically, we used a threshold of 0.2 for perspective and 0.1 for elevation. We validate these choices of thresholds by evaluating the correlation between the original metrics and converted accuracies in Figs. A14 and A15.

### v3

@@ 1 Introduction
Large-scale pre-trained vision models like Stable Diffusion [63] and DINOv2 [57], trained solely on 2D images, have shown strong generalization when fine-tuned on various computer vision tasks, including multi-view [39, 10] and monocular depth estimation [18, 23, 84], even reaching human-level performance in ordering objects by depth [49]. As a result, numerous recent monocular depth estimation approaches [29, 41, 80] leverage these models’ depth priors to achieve state-of-the-art performance. This naturally raises an important question: How does depth perception arise in these large vision models?
@@ 2 Related Work
Monocular Depth Estimation. Monocular depth estimation is an inherently ill-posed and challenging problem, yet deep learning-based methods have made significant strides in addressing it. Initial progress was driven by supervised methods [17, 20, 6] that were trained on various labeled datasets [55, 24]. Later, self-supervised approaches advanced the field further by leveraging stereo [22, 26] or temporal consistency [88, 28, 27]. More recently, general purpose methods [61, 79, 80, 41] leveraging large vision backbone models pre-trained with self-supervised [57] or generative objectives [63] and then fine-tuned on very large-scale supervised depth datasets have been shown to result in strong monocular depth estimation performance. However, it remains unclear how depth perception develops in the pre-trained models without explicit depth supervision during pre-training. In this paper, we try to shed light on the emergence of these monocular depth cues used by the human visual system in large vision models.
@@ 3.4 Perspective
Task. This cue is closely tied to vanishing point [50], which is defined as the point on the image plane where 2D perspective projections of 3D parallel lines intersect. Therefore, we use the task of vanishing point estimation to investigate the understanding of perspective. While there exist different versions of the task that define different numbers of vanishing points [69, 82], we adopt a simplified version that focuses on a single dominant vanishing point. Given an image $I$, the objective is to estimate the $\mathbf{p}=(x,y)$ coordinates of the vanishing point in the image coordinate system.
@@ 4.2 Probing Features
Following [57, 84], we perform a hyper-parameter search on the layers of model $\phi$. For each task, we extract features from different layers of $\phi$ which is frozen, and train a probe on these features under the same setting. The best layer is then selected based on the validation performance, and the probing on this layer is repeated five times using different random seeds. We report the mean and standard deviation of the test performance.
@@ 5.1 Monocular Depth Cues in Vision Models
Additional Baselines. During the probing of certain cues, we explicitly use mask information to construct the input feature, $\mathbf{f}$. Although these masks are not directly used by the probes, they could theoretically be leveraged to solve tasks. To verify that any additional information from this step is insufficient to solve the task, we designed a trivial baseline in which model features are replaced with simple 2D coordinate maps i.e., $\phi(I)$ in Eqs. 1 and 2 becomes $C\in\mathbb{R}^{H\times W\times 2}$ where $C_{h,w}=[h,w]^{T}$. Moreover, for each cue, we include an end-to-end baseline, where we train ResNet18 [30] and ViT-B/16 [16] models from scratch in an end-to-end fashion and select the best. We concatenate the masks with the input images channel-wise, allowing us to assess how much of each task can be solved without relying on sophisticated pre-trained features. Finally, to confirm that the tasks are readily solvable by humans, we evaluate our own performance on a random sample from the test sets, and obtained an accuracy of 95%$\pm$1.48across tasks.
Additional Tasks. To extend our analysis, we also evaluate vision models for monocular depth estimation (on NYUv2 [65] and DIW [12]) and image classification (on ImageNet-1k [15]) using a linear probe, a standard approach in previous work [57, 81, 3] for probing these datasets. Following the previous works [6, 12, 18], For NYUv2, we report accuracy as the percentage of pixels where the ratio between predicted and ground-truth depth is less than 1.25. We report top-1 accuracy for ImageNet. More details on these tasks are provided in Sec. A.1, and we show the linear probing results of these models on NYUv2 in Fig. 3.
@@ 5.3 Discussion
Limitations. Our analysis focuses only on widely used pre-trained model weights but not directly on the impact of various pre-training objectives and their training datasets on the emergence of depth cues. While depth from ego-motion and scene motion are also important cues for humans, our focus is primarily on assessing monocular cues learned by large vision models trained on static image collections and not image sequences (except CroCo, LRM, and DUSt3R).
@@ 6 Conclusions
Acknowledgments. Funding was provided by ELIAI (the Edinburgh Laboratory for Integrated Artificial Intelligence) - EPSRC (EP/W002876/1). HB was supported by the EPSRC Visual AI grant EP/T028572/1, and CJ was supported by a gift from Adobe.
@@ A.1 Full Benchmark Results
The probing results of all models are summarized in Tab. A1, and the average performance of models on DepthCues is plotted against their NYUv2 depth estimation (Fig. A1) as well as DIW depth ordering (Fig. A2) results. It is noted that there are slight differences between the linear probing results on NYUv2 and ImageNet-1k compared to previous works [57, 18]. These may stem from differences in the implementation details including the features probed, learning rates, and training iterations, etc. For example, the probing results on NYUv2 depth estimation in [18] are generally higher than our results. Possible reasons for this include the following factors. Firstly, while we train a simple linear layer to probe the patch tokens from the last layer of the vision models, a more complex non-linear probe similar to the DPT decoder [62] is used to probe features from multiple layers of the models in [18]. Secondly, the class tokens are utilized for several models (e.g., DINO, DINOv2) in [18], while we only use patch tokens for consistent evaluation across all models as some of the models do not include a class token. Thirdly, because of the difference in the probes, their optimization hyper-parameters can differ. Similar reasons can be applied to the differences in our ImageNet-1k probing results from previous works. In particular, for CLIP, the class token seems to be crucial for its ImageNet classification ability. We made these design choices because our primary focus is to evaluate the awareness of human-like monocular depth cues in vision models in a fair and comparable manner, rather than optimizing their probing performance on external benchmarks.
Figure A2: Performance of vision models on DepthCues vs. DIW depth ordering. A strong correlation is observed between depth cue understanding and depth estimation.
@@ A.5 Learning Depth Cues
Model NYUv2 Acc. (%) $\uparrow$ DIW WHDR (%) $\downarrow$
DINOv2 87.81 (0.09) 11.95 (0.05)
concat(DINOv2, DINOv2+DC ) 88.43 (0.09) 11.66 (0.11)
CLIP 43.82 (0.04) 35.33 (0.09)
concat(CLIP, CLIP+DC ) 44.40 (0.18) 32.69 (0.98)
Table A4: Linear probing on downstream depth estimation with different random seeds. Fine-tuning on DepthCues benchmark is also repeated.
In Sec. 5.2 of the main paper, we presented linear probing results on NYUv2 and DIW, showing that fine-tuning on DepthCues improves models’ performance on depth estimation. Here we further investigate whether the fine-tuned model has improved understanding of the depth cues by comparing the probing results of the fine-tuned and pre-trained DINOv2 on DepthCues. Note we focus on probing the last layer since LoRA was only applied on that layer during fine-tuning. The results from Tab. A3 shows that the fine-tuned version outperforms the original DINOv2 on four cues, with a slight drop in performance on light-shadow and perspective. Notably, we see a 9.4% increase in accuracy on texture-grad, on which the original DINOv2 shows relatively weaker understanding (see Fig. 3 in the main paper). In addition, we repeated the fine-tuning (on DepthCues) and evaluation of DINOv2 and CLIP on NYUv2 and DIW with three different random seeds, and show the results in Tab. A4, where significant improvements are observed.
@@ A.6 Monocular Depth Estimation Models
Relative vs. Metric Depth Models. The DepthAnythingv2 model checkpoint we evaluated so far was pre-trained for relative depth estimation. Here we additionally evaluate the two publicly available metric versions of DepthAnythingv2 on DepthCues, which are fine-tuned from the relative version on indoor and outdoor datasets respectively. It is observed from Tab. A5 that the metric versions achieved lower performance than the relative one. As metric models are fine-tuned from the relative one on smaller datasets, the difference in performance could also be impacted by the training data. We leave further investigation to future work.
DAv2 ver. Ele. Lit-shd. Occ. Prsp. Size Txt.-grd.
relative 83.73 84.65 81.16 95.33 82.16 89.90
abs. (outdoor) 74.96 82.91 78.60 92.00 83.85 88.70
abs. (indoor) 79.96 82.44 79.67 94.67 83.43 89.90
Table A5: Probing results of relative vs. metric versions of DepthAnyv2 on DepthCues.
@@ C.2 Evaluation Metrics
While we use accuracy to evaluate performance on the binary classification tasks, as discussed in the main paper we apply thresholding to the Euclidean distance (for perspective) and horizon detection error [^2] This measures the maximum distance between the predicted and the ground-truth horizon lines, normalized by the image height. [77] (for elevation) to convert these metrics into accuracy. Specifically, we used a threshold of 0.2 for perspective and 0.1 for elevation. We validate these choices of thresholds by evaluating the correlation between the original metrics and converted accuracies in Figs. A15 and A16.